OpenStudy (anonymous):

find the slope of the line passing through the pair of points (1,3), (-5,4)

OpenStudy (anonymous):

change in y divided by change in x, that is 4-3/5-1 = 1/4

OpenStudy (dumbcow):

slope formula m = (y2-y1)/x2-x1)

OpenStudy (anonymous):

so confused bc i got what frederik got but my only options for answers are -6, -1/6 , -7/4, and 6

OpenStudy (dumbcow):

answer is -1/6 x1 = 1 y1 = 3 x2 =-5 y2=4 m = (4-3)/-5-1) = -1/6

OpenStudy (anonymous):

Sorry, didn't see that (-)5